OpenStudy (anonymous):

Solve the equation for x: sinx=cosx I tried dividing both sides by cosx and using one of the quotient identities but I got the wrong answer

OpenStudy (lucaz):

use the unit circle

OpenStudy (irishboy123):

if \(sinx=cosx\), we can conclude something about x 

OpenStudy (irishboy123):

\(\large \sin{x} = \frac{o}{h}\) \(\large \cos{x} = \frac{a}{h}\)

OpenStudy (irishboy123):



OpenStudy (anonymous):

I keep getting pi over 4 and the answer is pi/6 + 2pi/3 n (n=any term because there isn't a limit)